OpenStudy (anonymous):

Solve the system by substitution. -4.5x-2y=-12.5 3.25x-y=-0.75 Please help my step by step so I can really understand, I'm super confused!

hero (hero):

Isolate y in the second equation to get y = 3.25x + 0.75

hero (hero):

Then substitute the expression for y in place of y in the first equation: I'd also multiply the first equation by 10 to make things easier: -45x -20y = -125 -45x - 20(3.25x + 0.75) = -125

hero (hero):

I know you may be uncomfortable with decimals, but there is not always a way to avoid them completely.

hero (hero):

This is one of those cases.

OpenStudy (anonymous):

so multiply like this? (10)-45x- (10)-20y= (10)-125

OpenStudy (anonymous):

@Hero